Brian buys 1 pack of baseball cards to add to the 2 cards a friend gave him. Then his mother gives him 2 more packs as a special treat. Now he has as many as cards s Marcus who owns 1 pack plus 12 loose cards. How many cards are in each pack
Answer by ikleyn(52946)

.

Let x be the number of card in a pack (assumed to be the same for the packs in this problem).


Write an equation as you read the problem

    x + 2 + 2x = x + 12   cards.


Simplify and find x

    3x + 2 = x + 12

    3x - x = 12 - 2

       2x  =    10

        x  =    10/2 = 5.


ANSWER.  There are 5 cards in each pack in this problem.

Solved.

--------------------

Today's morning tasks are much more reasonable than the usual ones at this time of a day.

The impression is as if the sun of reason shone in the east.